# SWAA: Sliding Window Attention Adaptation for Efficient Long-Context LLMs Without Pretraining

### v4

@@ 4.1 Models
All models are served with vLLM in float16 precision using a batch size of 64, with our customized Flash-attention-2 kernel to support Keep First and FA Decoding. We use greedy decoding (temperature =0=0) for all evaluations. In preliminary experiments, we observed that vLLM yields slightly lower (about 1% to 5%) scores than HuggingFace Transformers due to precision-related discrepancies.
4.2 Evaluation Dataset
@@
SWA is identical to full attention when the context length is within the window size. Even if the model is fine-tuned, we can simply disable the LoRA adapters for short prompts to get the same response as the original model. Therefore, our evaluations focus on long-context benchmarks, as re-evaluating models on standard short-context benchmarks (e.g., MMLU Hendrycks et al. (2021), GPQA Rein et al. (2023)) is completely unnecessary.
@@
We select LongMemEval Wu et al. (2024), a benchmark consisting of various types of long-context QA tasks with moderate difficulty, as our main benchmark, although it is originally designed for agent memory system evaluation. Its context length is controllable by selecting a specific number of chat sessions to concatenate as the context from a pool of hundreds of sessions (a session contains the chat history between user and assistant within a day). To create a moderately difficult and discriminative evaluation, we construct LongMemEval_24k by sampling 10 sessions, resulting in 500 samples ranging from 16k to 32k with an average context length of 24k.
@@
We also experiment on LongBench-V2 Bai et al. (2024b) and Ruler Hsieh et al. (2024), which are most commonly used long-context benchmarks, for additional validation of generalizability. For LongBench V2, we retain only the samples whose context length is under 128k due to GPU memory limitations; thus, 311 of 500 samples are kept. For Ruler, we choose the Multi-Query task, which contains 500 samples, and control the context length to 128k (counted by Qwen3 tokenizer).
To judge the correctness for accuracy, we use LLM-as-judge with GPT-5-Mini OpenAI (2025) for LongMemEval and exact-match for LongBench-V2 and Ruler.
However, we find other LongBench-V2 is too difficult for 4B-level models, and Ruler’s task types are not diverse enough (see details in Appendix B), so our analysis will more focus on the clearer results of LongMemEval.
4.3 Training Details
For the fine-tuning dataset, we initially considered LongAlign Bai et al. (2024a), a widely used long-context fine-tuning dataset for adapt a regular-length model to long-context tasks. However, since its sample count ($\sim$10,000) is insufficient, we incorporate an additional 6,000 samples from Fusang-v1-long Pan (2024), a more comprehensive corpus of over 40,000 long-context samples that includes LongAlign as a subset.
@@ 4.3 Training Details
We perform SWA-aware fine-tuning using LoRA Hu et al. (2022). Unless otherwise noted, we use rank $r=16$ and $\alpha=128$, and apply LoRA only to the query, key, and value projection modules. We adopt this parameter-efficient setting because full-parameter fine-tuning often leads to overfitting and degradation of the model’s original capabilities in our preliminary experiments. We use a learning rate of 1e-4 with a cosine decay schedule. Models are fine-tuned for a single epoch on the sampled long-context dataset since we observe no meaningful gains from additional epochs (see Appendix E). Training of each SWAA configuration takes approximately 12 hours on an 8*H20 GPU server for Qwen3-4B and 30 hours for Qwen3-30B-A3B.
5 Experiment Results
5.1 SWA Adaptation Without Fine-tuning
We first investigate training-free SWA adaptation. Table 1 reports LongMemEval_24k accuracy for Qwen3-4B-Thinking ("think") and Qwen3-4B-Instruct ("non-think") under different combinations of SWAA components. In most settings, we use an aggressive 2k window to amplify the impact of SWA. The configurations are grouped by the number of methods applied (0, 1, 2, or 3 of Interleaving Layers, Keep First and FA Decode). Rows 0 (original model) and 1 (naive SWA) serve as upper and lower baselines, respectively. In the column "FA layers", the value records which layers use full attention, and [] means no layers. In the column "Keep first", the value is $k$ in Keep First $k$ Tokens. From the results of LongMemEval, we find that:
Naive SWA is not viable. For LongMemeval, naively replacing FA with a 2k sliding window attention (row 1) drops accuracy significantly to $3.2$ and $11.0$, respectively. Even with an 8k window (row 2), accuracy only recovers to $13.2$ and $19.8$, far below the FA baseline.
Individual methods provide gains but fail to close the performance gap. Each method—Keep First, FA Decode, or Interleaving Layers—improves over naive SWA (rows 3–6), yet each alone recovers only a small fraction of the FA gap and remains well below the baseline. In short, no single method is sufficient.
Combinations exhibit strong synergy. Recipes that combine multiple methods deliver large gains: (1) FA Decode + Keep First $k$ substantially improves over naive SWA (rows 7–9), recovering approximately half to two-thirds of the performance drop for the thinking model as $k$ increases. However, increasing $k$ from 100 to 1000 yields little improvement, indicating that $k$ does not need to be exceedingly large. (2) Interleaving Layers + FA Decode is markedly stronger (row 13), recovering most of the gap for the thinking model. (3) FA Decode + Interleaving Layers + Keep First $k$ perform best (rows 18). The thinking model recovers close to 90% of the FA gap even at 2k window.
CoT synergizes with FA Decode. Under recipes that include FA Decode, the thinking model consistently benefits more than the non-thinking model (rows 13 and 18), suggesting that CoT synergizes with FA Decode: preserving global attention at decoding time enables longer reasoning traces to capitalize on context processed by SWA, confirming our hypothesis in Section 3.1.
Sliding window size affects, but is not the decisive role. With FA Decode + Keep First $k$, accuracy improves as the window increases (rows 7, 14, 15), though benefits are modest until 8k. When added with interleaving FA layers, moving from a 2k to a 4k window is enough to close the remaining gap on the thinking model (row 22 matches the FA baseline), indicating that FA Decode and Interleaving are the dominant drivers, and larger windows mainly smooth residual error.
SWA/FA layer selection has large impacts. Selecting less layers for FA, e.g. only one fourth or eighth (row 19, 20), though more efficient, will significantly decrease the improvement brought by Interleaving Layers. More importantly, for Qwen3-4B, configuring odd-numbered layers with full attention is significantly better than even-numbered layers (row 11, 13, 16, 17). However, surprisingly, this result is reversed for Qwen3-30B-A3B and Llama3.1-8B-Instruct (row 10 and 12 in Table 4 and 5). This suggests that layer functionalities differ across model families and sizes, necessitating model-specific layer selection strategies, as discussed in Section 3.2.
For LongBench V2, due to the high difficulty, the performance is generally poor. Some scores are even below the random guessing baseline (25%). However, despite narrower performance gaps between configurations, the overall accuracy trend still aligns with that of LongMemEval, except for few outliers, which may be attributed to the inherent randomness of multiple-choice questions.
For Ruler, as it involves only the needle-in-a-haystack task, performance varies drastically between configurations. The accuracy of the thinking model is even always lower than the non-thinking one. And the accuracy is hardly improved until we let at least half layers use FA and enable FA decoding at the same time.
Therefore, we answer RQ1: adapting an FA LLM to SWA is feasible even without any training. But it requires specific combinations of at least 2 methods, which could be less efficient for inference.
5.2 SWA Adaptation With Fine-tuning
We next evaluate SWA-aware supervised fine-tuning, which is expected to provide higher improvement. Table 2 reports LongMemEval_24k accuracy after SFT under various SWAA configurations. The original full-attention model is also fine-tuned as a baseline (row 0). Since training is relatively time-consuming, we only test a representative subset of configurations. From the results of LongMemEval, our findings are as follows:
Fine-tuning substantially lifts all SWA configurations. Comparing all the fine-tuning results with non-training ones, it is clear that fine-tuning consistently provides great improvement. However, simply fine-tuning with naive SWA remains insufficient, only achieving 18.8% and 23.8% accuracy (row 1).
FA Decode and Interleaving Layers emerge as dominant components. After SFT, FA Decode and Interleaving Layers provide the largest gains. Using FA Decode alone (row 3) or using Interleaving Layers alone (row 4) both get high accuracy. And combining both (row 8) further boosts performance to $73.2$ (think) and $58.8$ (non-think), nearly matching the full-attention SFT baseline in row 0.
Keep First becomes optional rather than essential. Before fine-tuning, Keep First is crucial for stability under SWA. But after SFT, it only provides minor additional improvement. With FA Decode, adding $k=100$ (row 7) improves over $k=0$ (row 3) only 4.5%, and if further combining FA layers, it almost offers no improvements (row 11 and row 8).
Effect of sliding window size. Row 3 and 5 shows that increasing the window from 2k to 4k with FA Decode improves thinking-model accuracy from $57.9$ to $62.6$. This mirrors the non-SFT trend that larger windows help, but the dominant improvements still come from FA Decode and Interleaving Layers.
The improvement on LongBench-V2 is mostly consistent with that on LongMemEval. While for Ruler, SFT also brings improvements, but configuring odd-numbered layers with full attention has a decisive effect that is far greater than other methods. Notably, long reasoning has no effect or even a negative effect on Ruler, which may also be the reason for the little improvement of FA decoding.
Thus, regarding RQ2, we conclude that fine-tuning yields substantial performance restoration, provided we apply FA Decode, Interleaving Layers, or a combination thereof, while Keep First becomes optional. And the improvement brought by SFT under each configuration varies significantly, implying that a configuration optimal for the training-free setting is not necessarily optimal after SFT, and vice versa.
@@ 5.2 SWA Adaptation With Fine-tuning
5.3 Performance–efficiency Trade-offs and Recommended Recipes
Although integrating more methods can typically achieve higher accuracy, it introduces more overhead, indicating that the efficiency of each recipe must also be evaluated. To assess the performance-efficiency trade-off of different SWAA configurations, we evaluate time-to-first-token (TTFT), time-per-output-token (TPOT), total throughput, and average running time per request. Concretely, we benchmark Qwen3-4B-Thinking on a single H100 GPU using vLLM’s bench_serve utility Kwon et al. (2023) with random input data and 100 total requests. The prompt length and output length are set to 128k and 512 tokens, respectively, representing a typical long-context QA setting.
To visualize the performance-efficiency trade-off, Figure 2 plots each configuration’s accuracy on LongMemEval_24k Wu et al. (2024) against its average running time, while detailed TTFT, TPOT, and throughput statistics for each configuration are provided in Appendix D. We draw a line between the full-attention point and the naive-SWA point as a baseline curve: configurations above this line offer a better accuracy-latency balance intuitively. For configurations with nearly identical time costs, we display only the one with the highest accuracy. Since Keep First $k$ has negligible impact on runtime (Appendix D), all plotted configurations fix $k=10$.
@@ 5.3 Performance–efficiency Trade-offs and Recommended Recipes
From Figure 2, we observe that many configurations in Figure 2 achieve a clearly better performance-efficiency ratio than baselines. For example, after SFT, (1) enable either FA decoding or Interleaving Layers can enhance the inference speed by almost 100%, while maintaining nearly 90% of the original accuracy; (2) while combining them can achieve nearly 100% original accuracy maintenance and over 30% acceleration; (3) By controlling the number of layers using FA, efficiency and quality can be achieved between that of (1) and (2). And for the thinking model, more points lie above the baseline curve compared to non-thinking, indicating that CoT generally has a positive effect on improving the performance-efficiency ratio of SWAA.
Thus, we finally answer RQ3: many SWAA configurations all reach excellent performance-efficiency trade-off, but there is no single metric to quantify such trade-off to decide the globally optimal one. We therefore summarize recommended SWA adaptation recipes tailored to various deployment scenarios in Table 3. And we must note that specific parameters should be flexibly set to meet application-specific requirements, without the need to follow our experimental parameters (e.g., a 2k window, $k=10$). For example, you can increase the window size to 4k or $k$ to 128 for higher accuracy and acceptable additional overhead.
@@ 6 Conclusion
In this work, we validate the feasibility of adapting full-attention pretrained LLMs to Sliding Window Attention (SWA) for better efficiency, offering a cost-effective alternative, SWAA, that avoids training sparse-attention models from scratch. By systematically deconstructing the adaptation process, we identify that the catastrophic degradation observed in naive implementations can be effectively mitigated through synergistic combinations of auxiliary methods. Our extensive experiments across Qwen and Llama families demonstrate that while trade-offs between computational overhead and model performance are inevitable, optimized configurations can achieve excellent performance-efficiency balance to effectively accelerate LLM inference.
7 Limitations
We hypothesize that the ideal reasoning trajectory of the model adapted to SWA should be longer than the original model, to compensate for the information loss caused by SWA. That means, using the answers generated by the original model as fine-tuning data may not be the optimal training method. Rather, RL methods like GRPO Shao et al. (2024) might further help the model adapted to SWA by learning a more adaptive reasoning trajectory. However, we did not experiment with them since they are too time-consuming and unstable.
We have not yet implemented the KV cache eviction (or overwriting) mechanism when using SWA; that is, although the speed is improved, the memory usage is actually unchanged. That is because not all of our SWAA configurations support discarding early KV cache.

### v5

@@ 4.1.1 Models
Our primary experiments use Qwen3-4B-Thinking and Qwen3-4B-Instruct (Team, 2025b). The Thinking variant enforces chain of thought (CoT) style reasoning, whereas the Instruct variant usually just answers briefly, so we can clearly see the impact of CoT on overcoming SWA’s structural defects. To ensure generality, we additionally evaluate Qwen3-30B-A3B-Thinking, Qwen3-30B-A3B-Instruct (Team, 2025b), and Llama3.1-8B-Instruct (Touvron et al., 2023), as shown in Appendix C. All models are served with vLLM in float16 precision using a batch size of 64, with our customized Flash-attention-2 kernel to support Keep First and FA Decoding. We use greedy decoding (temperature =0=0) for all evaluations. In preliminary experiments, we observed that vLLM yields slightly lower (about 1% to 5%) scores than HuggingFace Transformers due to precision related discrepancies.
4.1.2 Evaluation Dataset
@@
SWA is identical to full attention when the context length is within the window size. Even if the model is fine tuned, we can simply disable the LoRA adapters for short prompts to get the same response as the original model. Therefore, our evaluations focus on long context benchmarks, as re-evaluating models on standard short context benchmarks (e.g., MMLU (Hendrycks et al., 2021), GPQA (Rein et al., 2023)) is completely unnecessary.
@@
We select LongMemEval (Wu et al., 2024), a benchmark consisting of various types of long context QA tasks with moderate difficulty, as our main benchmark, although it is originally designed for agent memory system evaluation. Its context length is controllable by selecting a specific number of chat sessions to concatenate as the context from a pool of hundreds of sessions (a session contains the chat history between user and assistant within a day). To create a moderately difficult and discriminative evaluation, we construct LongMemEval_24k by sampling 10 sessions, resulting in 500 samples ranging from 16k to 32k with an average context length of 24k.
@@
We also experiment on LongBench-V2 (Bai et al., 2024b) and Ruler (Hsieh et al., 2024), which are most commonly used long context benchmarks, for additional validation of generalizability. For LongBench V2, we retain only the samples whose context length is under 128k due to GPU memory limitations; thus, 311 of 500 samples are kept. For Ruler, we choose the Multi-Query task (the most challenging needle-in-a-haystack task), which contains 500 samples, and control the context length to 128k (counted by Qwen3 tokenizer). To judge the correctness for accuracy, we use LLM as judge with GPT-5-Mini (OpenAI, 2025) for LongMemEval and exact match for LongBench-V2 and Ruler. However, we find other LongBench-V2 is too difficult for 4B level models, and Ruler’s task types are not diverse enough (see details in Appendix B), so our analysis will more focus on the clearer results of LongMemEval.
4.1.3 Training Details
@@
For the fine tuning dataset, we initially considered LongAlign (Bai et al., 2024a), a widely used long context fine tuning dataset for adapt a regular length model to long context tasks. However, since its sample count ($\sim$10,000) is insufficient, we incorporate an additional 6,000 samples from Fusang-v1-long (Pan, 2024), a more comprehensive corpus of over 40,000 long context samples that includes LongAlign as a subset. We perform SWA aware fine tuning using LoRA (Hu et al., 2022). Unless otherwise noted, we use rank $r=16$ and $\alpha=128$, and apply LoRA only to the query, key, and value projection modules. We adopt this parameter efficient setting because full parameter fine tuning often leads to overfitting and degradation of the model’s original capabilities in our preliminary experiments. We use a learning rate of 1e-4 with a cosine decay schedule. Models are fine tuned for a single epoch on the sampled long context dataset since we observe no meaningful gains from additional epochs (see Appendix E). Training of each SWAA configuration takes approximately 12 hours on an 8*H20 GPU server for Qwen3-4B and 30 hours for Qwen3-30B-A3B.
Table 1: The accuracy of Qwen3-4B on LongMemEval, LongBench V2, and Ruler with different SWAA recipes.
No. SFT Window FA layers Keep FA LongMem LB V2 Ruler
size first decode Think Non T Think Non T Think Non T
0 False Full / / / 73.0 62.0 34.6 35.2 85.6 92.8
1 False 2k [] 0 False 3.2 11.0 9.4 25.8 0.0 0.0
2 False 8k [] 0 False 13.2 19.8 15.1 22.1 0.0 0.0
3 False 2k [] 10 False 16.0 15.6 7.7 25.8 0.0 0.0
4 False 2k [] 0 True 11.8 14.2 26.2 25.2 0.2 0.0
5 False 2k [1, 3, 5, …] 0 False 13.4 18.4 12.1 23.5 0.0 0.0
6 False 8k [] 0 True 26.2 25.0 22.8 25.5 0.0 0.0
7 False 2k [] 10 True 38.2 20.6 25.8 25.2 1.0 0.0
8 False 2k [] 100 True 50.0 17.8 24.2 26.5 3.2 0.4
9 False 2k [] 1000 True 50.0 20.2 23.8 25.2 3.2 0.2
10 False 2k [0, 2, 4, …] 10 False 17.0 14.8 19.8 29.2 0.0 0.0
11 False 2k [0, 2, 4, …] 0 True 32.2 26.0 23.8 29.9 1.0 2.0
12 False 2k [1, 3, 5, …] 10 False 25.8 36.4 21.1 29.9 0.0 0.0
13 False 2k [1, 3, 5, …] 0 True 59.2 34.8 28.5 26.8 22.8 32.8
14 False 4k [] 10 True 38.0 24.4 27.9 27.5 1.6 2.2
15 False 8k [] 10 True 49.2 35.2 36.7 30.2 3.2 6.2
16 False 2k [0, 2, 4, …] 10 True 36.0 17.2 30.5 30.2 10.0 3.0
17 False 2k [1, 3, 5, …] 10 True 65.0 53.6 35.4 28.6 59.2 75.8
18 False 2k [1, 3, 5, …] 100 True 68.8 50.6 35.6 31.8 58.0 78.8
19 False 2k [1, 5, 9, …] 10 True 53.2 31.4 29.6 26.4 9.8 26.4
20 False 2k [1, 9, 17, …] 10 True 36.4 18.8 28.6 25.4 0.6 0.2
21 False 2k [3, 7, 11, …] 10 True 54.2 34.6 27.7 26.0 1.6 4.2
22 False 4k [1, 3, 5, …] 100 True 73.0 54.2 33.4 31.2 65.4 83.6
23 False 8k [1, 3, 5, …] 100 True 71.6 56.6 35.4 31.5 67.2 86.0
24 True Full / / / 74.6 63.4 37.9 34.9 88.2 90.6
25 True 2k [] 0 False 18.8 23.8 7.4 30.9 0.0 0.0
26 True 2k [] 100 False 15.6 24.0 6.0 30.2 0.0 0.0
27 True 2k [] 0 True 57.9 42.0 29.2 30.2 0.4 0.8
28 True 2k [1, 3, 5, …] 0 False 63.6 54.6 29.5 31.9 59.0 62.2
29 True 2k [] 10 True 59.0 41.4 28.9 29.2 9.2 14.4
30 True 2k [] 100 True 62.2 42.6 29.2 30.5 13.4 20.4
31 True 2k [1, 3, 5, …] 0 True 73.2 58.8 38.3 34.6 74.0 85.8
32 True 2k [0, 2, 4, …] 0 True 66.0 $\backslash$ 31.5 $\backslash$ 8.0 $\backslash$
33 True 2k [1, 5, 9, …] 0 True 68.8 47.0 32.0 32.2 7.6 28.0
34 True 2k [1, 3, 5, …] 100 True 73.2 61.4 37.2 33.9 72.8 89.2
4.2 Experiment Results
4.2.1 Accuracy Analysis
A macroscopic analysis of our experimental results of accuracy (Table 1) reveals a fundamental principle: the catastrophic degradation caused by SWA cannot be resolved by addressing the dual bottlenecks, training inference distribution drift and structural severing of distant semantic flow, in isolation. Applying any single method, such as Keep First, FA Decode, Interleaving Layers, or naive SFT, yields negligible accuracy improvements over pure SWA (Table 1, rows 3 to 5 and 25). Conversely, when we pair methods targeting distinct challenges (e.g., Fine tuning and FA Decode, Fine tuning and Interleaving layers, or Keep First and FA Decode), we observe significant accuracy recoveries.
FA Decode and Temporal Retrieval Synergy. Applying SWA strictly during prefilling while reverting to FA during decoding provides a temporal selection mechanism that temporarily reopens direct pathways to distant tokens. We find this structural mitigation exhibits a profound synergy with the use of CoT, i.e. thinking. Because CoT naturally extends the auto regressive generation phase, it provides the model with more sequential steps to actively retrieve and integrate the broad context that was structurally under processed during the prefilling stage. This temporal retrieval synergy is validated by the superior performance of the thinking model over the standard instruct variant under FA Decode configurations (rows 13 and 18).
Interleaving FA and SWA Layers: Spatial Information Highways. While FA Decode addresses temporal bottlenecks, interleaving FA layers establishes vital spatial pathways, allowing distant information to propagate through the network’s depth. This is a highly potent structural mitigation, but zero shot adaptation is sensitive to layer selection. For instance, configuring odd numbered layers with FA vastly outperforms even numbered ones for Qwen3-4B (row 13 vs. row 11). This asymmetry suggests that specific layers in pretrained LLMs inherently function as global routing hubs, and misaligning SWA with these hubs disrupts the semantic flow.
Synergy of Structural Methods. Interestingly, combining two purely structural interventions, FA Decode and Interleaving Layers, also delivers substantial gains (row 13). This occurs because structural improvements selectively enable SWA, naturally nudging the overall attention pattern closer to its original full attention distribution. Consequently, this implicitly alleviates the training and inference mismatch alongside resolving structural deficits.
Keep First $k$ Tokens: Training free Attention stabilizer. Models pretrained with Full Attention (FA) develop a strong reliance on initial “sink” tokens. Prior to fine tuning, naively applying SWA cuts off access to these sinks, causing severe attention distribution drift. Preserving attention to these first $k$ tokens acts as a crucial distributional anchor, preventing complete generation collapse (rows 7 and 9). However, its effect is obviously weaker than, and will be covered by SFT, which intrinsically reshape the attention weights to not relying heavily on these sinks, rendering the Keep First strategy largely optional (row 27 vs. 29).
Fine Tuning: Aligning Model Weights with Sparse Attention Patterns. While training free methods provide partial recovery, SFT acts as the ultimate unifier, parametrically aligning the model’s internal representations with the modified sparse attention structures. Combining structural methods (FA Decode or Interleaving Layers) with lightweight fine tuning virtually eliminates the performance gap, restoring accuracy to perfectly match the full attention baseline (rows 27, 28, and 31).
The above findings are mainly based on the results of LongMemEval and LongBench V2, which are more general. While for Ruler, as it involves only the needle in a haystack task, it exhibits a more extreme pattern: the accuracy can hardly be improved unless we enable Interleaving Layers, coupled with at least one another method (row 13 and 28). This indicates the retrieval behavior may be concentrated in certain specific layers, therefore these layers must retain full attention.
@@ 4.2.1 Accuracy Analysis
4.2.2 Efficiency Analysis and Recommended Recipes
Although every method is helpful to accuracy recovery, Interleaving Layers and FA Decode introduce varying degrees of computation overhead in inference compared to pure SWA, indicating that the performance-efficiency trade-off of each SWAA configurations must also be evaluated, to find the most cost-effective one. So, we evaluate the average running time per request of Qwen3-4B-Thinking on a single H100 GPU using vLLM’s bench_serve utility (Kwon et al., 2023) with random input data and a total of 100 requests. The prompt length and output length are set to 128k and 512 tokens, respectively, representing a typical long context QA setting. More specific metrics such as time to first token (TTFT), time per output token (TPOT) and total throughput are shown in Appendix D.
To visualize the performance efficiency trade off, Figure 2 plots each configuration’s accuracy on LongMemEval_24k (Wu et al., 2024) against its average running time, while detailed TTFT, TPOT, and throughput statistics for each configuration are provided in Appendix D. We draw a line between the full attention point and the naive SWA point as a baseline curve: configurations above this line offer a better accuracy latency balance intuitively. For configurations with nearly identical time costs, we display only the one with the highest accuracy. Since Keep First $k$ has negligible impact on runtime (Appendix D), all plotted configurations fix $k=10$.
@@ 4.2.2 Efficiency Analysis and Recommended Recipes
From Figure 2, we observe that many configurations achieve a clearly better performance efficiency ratio than baselines. For example, after SFT, (1) enabling either FA decoding or Interleaving Layers can enhance the inference speed by almost 100%, while maintaining nearly 90% of the original accuracy; (2) while combining them can achieve nearly 100% original accuracy maintenance and over 30% acceleration; (3) By controlling the number of layers using FA, efficiency and quality can be achieved between that of (1) and (2). Furthermore, for the thinking model, more points lie above the baseline curve compared to the non thinking model, indicating that CoT generally has a positive effect on improving the performance efficiency ratio of SWAA.
Thus, we conclude that many SWAA configurations reach excellent performance efficiency trade offs, but there is no single metric to quantify such trade offs to decide the globally optimal one. We therefore summarize recommended SWA adaptation recipes tailored to various deployment scenarios in Table 2. We must note that specific parameters should be flexibly set to meet application specific requirements, without the need to follow our experimental parameters (e.g., a 2k window, $k=10$). For example, users can increase the window size to 4k or $k$ to 128 for higher accuracy and acceptable additional overhead.
@@ 5 Conclusion
In this work, we validate the feasibility of adapting full attention pretrained LLMs to Sliding Window Attention (SWA) for better efficiency, offering a cost effective alternative, SWAA, that avoids training sparse attention models from scratch. By systematically deconstructing the adaptation process, we identify that the catastrophic degradation observed in naive implementations can be effectively mitigated through synergistic combinations of auxiliary methods. Our extensive experiments across Qwen and Llama families demonstrate that while trade offs between computational overhead and model performance are inevitable, optimized configurations can achieve excellent performance efficiency balance to effectively accelerate LLM inference.